# Propagation of Interplanetary Shocks in the Heliosphere [^1] July 17, 2024

### v1

@@ 7 Summary and conclusions
In this paper, two conjugated multi-spacecraft studies of interplanetary (IP) shock propagation are presented. It aims to determine how IP shocks develop and evolve in spatial and temporal propagation. The spatial separation of observations is 40 million km. In this scale, we do not see any significant development of the IP shocks. The shock parameters are similar within their estimated error at each location. No sign of wave dispersion was observed.
The Event May 7, 2007 is a fast-forward (FF) shock event, appears planar and tilted $56.42^{\circ}$ to the Sun-Earth line. The tilt of this planar shock surface is almost identical to the usual propagation of the Parker spiral impacting the Earth. Therefore, the origin of the shock is co-rotating stream interactions (CIRs), which agrees with the detected CIR on that day. There is no sign of temporal change in this scale. In the spatial range of 40 million km, no change was observable. The G1-minor geomagnetic storm happened after this shock was detected, indicating this shock event caused the geomagnetic storm on that day.
The Event April 23, 2007 is a fast reverse (FR) shock event and, it appears the shape of the shock is not uniform and twisted from the STEREO--A spacecraft to the other spacecraft along the transverse direction to the Sun-Earth line. The shock normal vectors also were observed to be changing along the Y-axis. A G1-minor geomagnetic storm occurred on this day, similar to the fast-forward shock event of May 7, 2007. Intriguingly, the onset of this storm took place three hours before the shock detection time recorded by STEREO--A. The termination of the storm partially coincided with the STEREO--A detection time. The source of this shock is SIR/CIR. The G1-minor geomagnetic storm, with a $K_{p}$-5 index, began before the shock detection. It seems that the fast-moving solar wind instigated the minor storm before the arrival of the fast-reverse shock at the spacecraft’s position. The orientation irregularity can be potentially accounted for by the characteristic tilts of SIRs.

### v2

@@ 7 Summary and conclusions
In this paper, two conjugated multi-spacecraft studies of interplanetary (IP) shock propagation are presented. It aims to determine how IP shocks develop and evolve in spatial and temporal propagation. The spatial separation of observations is around 40 million km. In this scale, we do not see any significant development of the IP shocks. The shock parameters are similar within their estimated error at each location. No sign of wave dispersion was observed.
The Event May 7, 2007, is a fast-forward (FF) shock event, appears planar and tilted $56.42^{\circ}$ to the Sun-Earth line. The tilt of this planar shock surface is almost identical to the usual propagation of the Parker spiral impacting the Earth. Therefore, the origin of the shock is corotating interaction region (CIR), which agree with the detected CIR on that day. There is no sign of temporal change in this scale. In the spatial range of 40 million km, no change was observable. The G1-minor geomagnetic storm happened after this shock was detected, indicating this shock event caused the geomagnetic storm on that day.
The Event of April 23, 2007 is a fast reverse (FR) shock event and, it appears that the shape of the shock is not uniform and twisted from the STEREO--A spacecraft to the other spacecraft along the transverse direction to the Sun-Earth line. The shock normal vectors also were observed to be changing along the Y-axis. A G1-minor geomagnetic storm occurred on this day, similar to the fast-forward shock event of May 7, 2007. Intriguingly, the onset of this storm took place three hours before the shock detection time recorded by STEREO--A. The termination of the storm partially coincided with the STEREO--A detection time. The source of this shock is SIR/CIR. The G1-minor geomagnetic storm, with a $K_{p}$-5 index, began before the shock detection. It seems that the fast-moving solar wind instigated the minor storm before the arrival of the fast-reverse shock at the spacecraft’s position. The orientation irregularity can be potentially accounted for by the characteristic tilts of SIRs.